Suppose a gold miner finds a gold nugget and sells the nugget to a mining company for $500. The mining company melts down the gold, purifies it, and sells it to a jewelry maker for $1,000. The jewelry maker fashions the gold into a necklace that it sells to a department store for $1,600. Finally, the department store sells the necklace to a customer for $2,100. instructions: Round your answer to the nearest dollar As a result of these transactions, GDP increases by
Answer #1

GDP increases by the amount of value of added = (500 - 0) + (1000 - 500) + (1600 - 1000) + (2100 - 1600)

= 500 + 500 + 600 + 500 = 2100

So, GDP increases by $ 2100.